# IRIF, Université Paris Cité, France

### v1

@@ 3.2 Adversarial Resources and Goal
  • Known architecture. The attacker knows the architecture of the target neural network $f$, meaning she knows the depth of $f$ and the width of each layer.
@@
Adversarial goal. The objective of the extraction is not to exactly replicate the target network, but rather to achieve what is known as an $(\epsilon,\delta)$-functionally equivalent extraction [8]. We slightly change the definition to normalise the difference between the target and the extracted network.
@@
3.3 On Functionally Equivalent Extraction
We note four key differences between the extracted and target networks and justify that they will not affect the $(\epsilon,\delta)$-assessment.
Permutation of neurons. During the attack, we will extract a neuron from a hidden layer, but we will not know to which row (neuron in the layer) it corresponds. For example, if we denote the first extracted neuron as $\hat{A}_{1}^{(i)}$, we do not know if $\hat{A}_{1}^{(i)}=A_{1}^{(i)}$ or if $\hat{A}_{1}^{(i)}=A_{2}^{(i)}$, etc. We are therefore permuting the rows of $A^{(i)}$. However, when extracting the next hidden layer $\hat{A}^{(i+1)}$, the process naturally permutes the columns to match the permuted rows of $\hat{A}^{(i)}$, since $\hat{A}^{(i)}$ (rather than $A^{(i)}$) is used in the extraction. Therefore, this permutation does not compromise the overall extraction process over the layers. For the last output layer, we must directly align our extracted matrix with the output of $f$ as this ensures that this last recovered layer has the correct row order. We can conclude that the permutation of extracted neurons in a hidden layer affects neither $\epsilon$ nor $\delta$.
Scaling of neurons. When extracting a neuron $\eta^{(i)}_{k}$ (row $k$ from $A^{(i)}$), its values are scaled by a constant $c^{(i)}_{k}$. This scaled row is referred to as the signature of $A^{(i)}_{k}$. Since the extraction is performed on $\hat{A}^{(i)}$ rather than on $A^{(i)}$, every corresponding column in $A^{(i+1)}$ must be adjusted by dividing its values by $c^{(i)}_{k}$ to match $\hat{A}^{(i)}$. Even though $\hat{a}^{(i+1)}_{j,k}=\frac{a^{(i+1)}_{j,k}}{c^{(i)}_{k}}\cdot c^{(i+1)}_{j}$ for any $a^{(i+1)}_{j,k}\in A^{(i+1)}$, we still have,
$\hat{\Gamma}^{(i+1)}=\begin{pmatrix}c^{(i+1)}_{1}&0&\cdots&0&\\ 0&c^{(i+1)}_{2}&\cdots&0&\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&c_{d_{i+1}}^{(i+1)}\end{pmatrix}\cdot\Gamma^{(i+1)}.$
Once again, because the last output layer must match the output of the network, we have $\hat{a}^{(r+1)}_{j,k}=\frac{a^{(r+1)}_{j,k}}{c^{(r)}_{k}}\cdot 1$ and therefore $\hat{\Gamma}=\hat{\Gamma}^{(r+1)}=\Gamma^{(r+1)}=\Gamma$. Thus, here again, scaling does not affect $\epsilon$ and $\delta$.
Always-on/off neurons. Some neurons do not have critical points within the space of queries the attacker will make. They are either always active (on) or always inactive (off) in this input space. The attacker relies on critical points to identify hyperplanes and detect the existence of neurons. As such, always-on and always-off neurons cannot be detected or recovered through extraction. Always-off neurons can safely be ignored as they do not contribute to the network. Always-on neurons do not induce additional critical hyperplanes in $\mathbb{R}^{d_{0}}$ and make a linear map $f$ on the input space. The extraction naturally integrates this map when recovering the last layer of $f$. The authors of [6] reported that no always-on neurons were found on trained neural networks, but they did find instances of always-off neurons. In some cases, it can be relevant to extract the network on a small interesting subset $S$ of $\mathbb{R}^{d_{0}}$, where some neurons might be always-on. For example, if the target is a neural network tasked with image recognition, such a subset can be the hyperplane corresponding to black-and-white pictures. Extracting the neural network only on this hyperplane will make the attack less costly in time and in queries.
Dead weights. In some cases, some weights can be set to 00 without affecting the effectiveness of the extracted network $\hat{f}$. We refer to these weights as dead weights. A key instance comes from always-off and always-on neurons when we assume the correct extraction of all previous layers. Specifically, if $\eta_{k}^{(i)}$ is always on or always off, the weights on the $k$-th column of $A^{(i+1)}$ are dead weights as their contribution is absorbed by the last layer as explained above. We will come back to them in Section 5.
3.4 Overview of the Existing Attack
In this section, we recall the signature extraction algorithm from  [8]. The aim, as discussed earlier, is to extract an $(\epsilon,\delta)$-functionally equivalent network. The adversarial goals and the resources required for the extraction are outlined in Section 3.2. The extraction is performed layer by layer: first, we recover the parameters of layer 1, then use them to reconstruct layer 2, and so on. We now explain the process of recovering layer $i$, assuming that the first $i-1$ layers of the target network $f$ have been correctly extracted.
The recovery of each layer consists of two parts. The first part is a signature recovery (following [8]), which recovers each matrix row $A^{(i)}_{k}$ up to a scaling factor. The second part is sign recovery (following [13]’s tweak of [6]) which determines the sign of the scaling factor for each neuron. As our focus is on signature recovery, we refer the reader to  [13] and [6] for a description of the sign recovery.
Signature recovery is carried out in five main steps: searching critical points, recovering partial signatures, merging partial signatures, finding missing entries in the signatures, and computing the bias.
@@ 3.4 Overview of the Existing Attack
Partial signature recovery. When we cross a neuron’s critical hyperplane, we move, depending on the sign of the neuron, from a region (polytope) where the neuron does not contribute to the network’s output to one where it does. All other neurons remain unchanged. Therefore, by making queries close to a critical point, we can construct a system of equations that isolates the contribution of the neuron in question, allowing us to recover its parameters up to a sign. We now describe this process in more detail.
We define the second-order differential operator of $f$ along direction $\Delta$ as $\partial^{2}_{\Delta}f(x):=f(x+\epsilon\Delta)+f(x-\epsilon\Delta)-2f(x)$. Suppose the target layer is layer $i$. We assume that the previous layers have been extracted and thus that $\Gamma_{x}^{(i-1)}$ is known for all inputs $x$.
@@
$\partial^{2}_{\Delta}f(x)=c_{k}^{(i)}\big{|}(\Gamma_{x}^{(i-1)}\Delta)\cdot A_{k}^{(i)}\big{|}$
@@
$\frac{\partial^{2}_{\Delta}f(x)}{\partial^{2}_{\Delta_{0}}f(x)}=\frac{\big{|}(\Gamma_{x}^{(i-1)}\Delta)\cdot A_{k}^{(i)}\big{|}}{\big{|}(\Gamma_{x}^{(i-1)}\Delta_{0})\cdot A_{k}^{(i)}\big{|}}.$
@@
we can eliminate the absolute value (see Appendix 0.C.2), obtaining,
$\frac{(\Gamma_{x}^{(i-1)}\Delta)\cdot A_{k}^{(i)}}{(\Gamma_{x}^{(i-1)}\Delta_{0})\cdot A_{k}^{(i)}}.$
By taking enough directions $\Delta$ in the input space, we can solve for $A_{k}^{(i)}$ up to a constant, reconstructing the signature (see Appendix 0.C.3). Since $\Gamma_{x}^{(i-1)}=I_{x}^{(i-1)}A^{(i-1)}\Gamma_{x}^{(i-2)}$, ReLUs on layer $i-1$ are blocking some coefficients for all directions $\Delta$ we take around $x$. Therefore, instead of extracting a full (scaled) matrix row, for example $\left(a_{1},a_{2},a_{3},a_{4},a_{5}\right)$, we can only retrieve a partial signature, say, $\left(0,a_{2},a_{3},0,a_{5}\right)$.
Merging signatures into components. To reconstruct the full signature, we need to merge partial signatures obtained from different critical points of the same neuron $\eta_{k}^{(i)}$. Each recovered signature from critical points of $\eta_{k}^{(i)}$ is $A_{k}^{(i)}$ with some missing entries, scaled to an unknown factor. Assuming that no two rows of the matrix are identical, merging two signatures requires only checking whether all their shared non-zero weights are proportional. For example, consider two signatures $(\lambda a,\lambda b,0,\lambda d,0)$ and $(\Lambda a,\Lambda b,0,0,\Lambda e)$ where $\lambda$ and $\Lambda$ are constants. If $\frac{\lambda a}{\lambda b}=\frac{\Lambda a}{\Lambda b}$, they can be merged to get $(a,b,0,d,e)$, up to a constant (see Figure 2). The resulting merged row is called a component. The size of a component refers to the number of critical points whose partial signatures have merged to get the component.
If a component has a size greater than 2, we consider it to belong to the target layer, as it is highly unlikely for signature extraction to yield consistent results when applied to critical points from deeper layers. We repeat the process described above — randomly searching for critical points, extracting partial rows, and merging signatures — until we obtain at least $d_{i}$ components of size greater than 2, or until a predefined threshold on the number of critical points is reached (to prevent infinite loops caused by always-off or always-on neurons). If the number of components exceeds $d_{i}$, the number of neurons in layer $i$, we retain only the largest ones.
@@
Figure 2: Network where neurons on the previous layer are labelled in grey on their active side. In red is the neuron we aim to find. The critical points on the left and right yield respectively $(\Lambda a,\Lambda b,0,0,\Lambda e)$ and $(\lambda a,\lambda b,0,\lambda d,0)$. We can infer $(a,b,0,d,e)$ up to a constant, despite the hyperplane in red never entering a polytope where $\eta_{1},\eta_{2},\eta_{4}$ and $\eta_{5}$ are active.
@@
Recovering the bias. Once we have recovered the signature of $\eta^{(i)}_{k}$, $c^{(i)}_{k}A^{(i)}_{k}$, it suffices to pick a critical point $x$ of the component to compute the corresponding scaled bias $c^{(i)}_{k}b^{(i)}_{k}$ using the equation
$(c^{(i)}_{k}A^{(i)}_{k})\cdot F^{(i-1)}(x)+c^{(i)}_{k}b^{(i)}_{k}=0.$
@@
Last layer recovery. The oracle returns the complete raw outputs of $f$ and the last output layer is a linear layer. Therefore its extraction is straightforward. No additional query is needed for extraction, as we can reuse previously queried critical points to build the linear system.
4 Improving the Signature Recovery
To the best of our knowledge, no prior work has successfully applied signature extraction on deep enough neural networks. Specifically, according to Table 1 in [8], the deepest model the authors of this paper attempted to attack had an architecture of $40-20-10-10-1$, consisting of only $3$ hidden layers. On the other hand, Table 2 in [13] shows that they attempted to attack a model with an architecture of $1784-16^{(8)}-$ 1, which had $8$ hidden layers. However, despite running the attack for over $36$ hours, they failed to recover the fourth hidden layer. Although [6] demonstrated that their sign recovery was effective across all layers of a complex $13072-256^{(8)}-$ 1 model, they did not conduct experiments on signature recovery.
Motivated by these limitations, we analyze the signature extraction method and identify two challenges preventing its successful application on deeper layers. We then propose solutions to overcome them. The left part of Figure 3 illustrates the signature extraction workflow and highlights where these challenges arise.
We provide a brief overview of each failure. First, the signature extraction relies on solving a system of equations to recover the weights of the target neuron associated with all active neurons in the previous layer. However, as we go deeper into the network, the rank of this system might not match the number of active neurons on the previous layer. This mismatch can lead to an underdetermined system, resulting in incorrect signature extraction. Second, the component selection of the largest $d_{i}$ components of size greater than 2 is incorrect. Our experiments show that this strategy is only valid for shallow networks with comparatively fewer critical points on deeper layers. Figure 3 gives an outline of our improvements.
@@ 4 Improving the Signature Recovery
Figure 3: Left: Original attack from [8]. Right: Proposed improvements. Two error-inducing steps in the original attack are coloured on the left. Improvements match the colour of the step they address.
4.1 Identifying Deeper Merges

### v2

@@ 2.2 Adversarial Resources and Goal
  • Known architecture. The attacker knows the architecture of the target neural network $f$.
@@
Adversarial goal. The objective of the extraction is not to exactly replicate the target network, but rather to achieve what is known as an $(\epsilon,\delta)$-functionally equivalent extraction [8]. We slightly change the original definition to normalise the difference between the target and the extracted network.
@@
2.3 Overview of the Existing Signature Extraction
In this section, we recall the signature-extraction algorithm from [8]. The extraction is performed layer by layer: first, we recover the parameters of layer 1, then use them to reconstruct layer 2, and so on. We now explain the process of recovering layer $i$, assuming that the first i1i-1 layers of the target network $f$ have been correctly extracted.
Signature extraction is carried out in five main steps: searching critical points, recovering partial signatures, merging partial signatures, finding missing entries in the signatures, and computing the bias.
@@ 2.3 Overview of the Existing Signature Extraction
Partial signature recovery. When we cross a neuron’s critical hyperplane, we move, depending on the sign of the neuron, from a region (polytope) where the neuron does not contribute to the network’s output to one where it does. All other neurons remain unchanged. Consequently, by querying the network close to a critical point, we can construct a system of equations that isolates the contribution of the neuron in question, allowing us to recover its parameters up to a sign. We now describe this process in more detail.
Following [8], we define the second-order differential operator of $f$ along direction $\Delta$ as
$\partial^{2}_{\Delta}f(x):=\frac{1}{\epsilon_{1}}(f(x+\epsilon_{2}\Delta)+f(x-\epsilon_{2}\Delta)-2f(x)).$
Suppose the target layer is layer $i$. We assume that the previous layers have been extracted and thus that $\Gamma_{x}^{(i-1)}$ is known for all inputs $x$.
@@
$\partial^{2}_{\Delta}f(x)=c_{k}^{(i)}\big|A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta)\big|$
@@
$\frac{\partial^{2}_{\Delta}f(x)}{\partial^{2}_{\Delta_{0}}f(x)}=\frac{\big|A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta)\big|}{\big|A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta_{0})\big|}.$
@@
we can eliminate the absolute value (see Appendix 0.C.2), obtaining
$\frac{A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta)}{A_{k}^{(i)}\cdot(\Gamma_{x}^{(i-1)}\Delta_{0})}.$
By taking enough directions $\Delta$ in the input space, we can solve for $A_{k}^{(i)}$ up to a constant $c^{(i)}_{k}$, reconstructing the signature (see Appendix 0.C.3). We note $\hat{A}_{k}^{(i)}$ the recovered signature of $A_{k}^{(i)}$. As $k$ remains unknown, the extraction permutes the rows of $A^{(i)}$: we do not know if $\hat{A}_{1}^{(i)}=A_{1}^{(i)}$ or if $\hat{A}_{1}^{(i)}=A_{2}^{(i)}$. When extracting the next hidden layer $\hat{A}^{(i+1)}$, the end-to-end attack naturally permutes the columns to match the permuted rows of $\hat{A}^{(i)}$, since $\hat{A}^{(i)}$, rather than $A^{(i)}$, is used in the extraction. In the same fashion, the next layer absorbs the constant $c^{(i)}_{k}$: $\hat{a}^{(i+1)}_{j,k}=a^{(i+1)}_{j,k}\cdot\frac{c^{(i+1)}_{j}}{c^{(i)}_{k}}$. The extracted last layer forces the permutation and scaling to match the target neural network, yielding a functionally equivalent extraction.
However, since $\Gamma_{x}^{(i-1)}=I_{x}^{(i-1)}A^{(i-1)}\Gamma_{x}^{(i-2)}$, ReLUs on layer i1i-1 are blocking some coefficients for all directions $\Delta$ we take around $x$. Therefore, instead of extracting a full (scaled) matrix row, for example $\left(a_{1},a_{2},a_{3},a_{4},a_{5}\right)$, we can only retrieve a partial signature, say, $\left(0,a_{2},a_{3},0,a_{5}\right)$.
Merging signatures into components. To reconstruct the full signature, we need to merge partial signatures obtained from different critical points of the same neuron $\eta_{k}^{(i)}$. Each recovered signature from critical points of $\eta_{k}^{(i)}$ is $A_{k}^{(i)}$ with some missing entries, scaled to an unknown factor. Assuming that no two rows of the matrix are identical, merging two signatures requires only checking whether all their shared non-zero weights are proportional. For example, consider two signatures $(\lambda a,\lambda b,0,\lambda d,0)$ and $(\Lambda a,\Lambda b,0,0,\Lambda e)$ where $\lambda$ and $\Lambda$ are constants. If $\frac{\lambda a}{\lambda b}=\frac{\Lambda a}{\Lambda b}$, they can be merged to get $(a,b,0,d,e)$, up to a constant (see Fig. 1). The resulting merged row is called a component. The size of a component is the number of critical points whose partial signatures merged to form that component.
@@
Figure 1: Input space of a network where neurons on the previous layer are labelled in grey on their active side. In red is the neuron we aim to find. The critical points on the left and right yield respectively $(\Lambda a,\Lambda b,0,0,\Lambda e)$ and $(\lambda a,\lambda b,0,\lambda d,0)$. We can infer $(a,b,0,d,e)$ up to a constant, even though no single polytope activates $\eta_{1},\eta_{2},\eta_{4},\eta_{5}$ simultaneously.
If a component has a size greater than 2, the original attack from Carlini et al. [8] assigns it to the target layer, based on the belief that signature extraction would not produce consistent results on critical points from deeper layers (deeper layers are not expected to have such linear dependencies). However, we will explain later why this is not a correct perception, as critical points from deeper layers might indeed be merged. The process described above (randomly searching for critical points, extracting partial rows, and merging signatures) is repeated until the number of components of size >2>2 is equal to $d_{i}$, the number of neurons at layer $i$.
@@
Recovering the bias. Once we have recovered the signature of $\eta^{(i)}_{k}$, $c^{(i)}_{k}A^{(i)}_{k}$, it suffices to pick a critical point $x$ of the component to compute the corresponding scaled bias $c^{(i)}_{k}b^{(i)}_{k}$ using the equation $(c^{(i)}_{k}A^{(i)}_{k})\cdot F^{(i-1)}(x)+c^{(i)}_{k}b^{(i)}_{k}=0$.
@@
Last layer recovery. The oracle returns the complete raw outputs of $f$ and the last output layer is a linear layer. Therefore, its extraction is straightforward. No additional query is needed for extraction, as we can reuse previously queried critical points to build the linear system and recover the last linear layer’s weights.
2.4 Overview of Existing Sign-Extraction Methods
After recovering each neuron’s signature in layer up to a constant $c^{(k)}_{i}$, we must determine the sign of each constant. Canales-Martínez et al. [6] propose two complementary sign-extraction techniques: the system-of-equations (SOE) approach and the neuron wiggle. We briefly summarize both methods.
SOE. The system-of-equations method aims at contractive layers, where it recovers the signs of every neuron in the target layer together. Locally around an input $x$, for a small perturbation direction $\Delta_{k}$, we have
$f(x+\Delta_{k})-f(x)\;=\;G^{(i+1)}_{x}\,I^{(i)}_{x}\,A^{(i)}\,F^{(i-1)}_{x}\,\Delta_{k},$
where $G^{(i+1)}_{x}$ is the linear contribution of deeper layers at $x$, $A^{(i)}$ is the weight matrix of the target layer, $I^{(i)}_{x}$ is the diagonal activation mask of layer $i$ at $x$, and $F^{(i-1)}_{x}$ is the local extracted linear map up to layer i1i-1. To find the signs of all neurons, we want to find $I^{(i)}_{x}$. Equivalently, we want to find which entries of $G^{(i+1)}_{x}I^{(i)}_{x}$ are zero. Therefore we build the following system:
$\{c\cdot y_{k}=z_{k}\}_{k}$
where $c=G^{(i+1)}_{x}I_{x}^{(i)}$ is the variable we are looking for, $y_{k}=A^{(i)}F^{(i-1)}_{x}\Delta_{k}$ and $z_{k}=f(x+\Delta_{k})-f(x)$. To solve SOE, we thus need that $\text{rank}(F^{(i-1)}_{x})\geq\text{rank}(A^{(i)})$. This technique can therefore be used as long as the network stays very contractive. It is thus usually applicable only for the first layer.
Neuron Wiggle. For non-contractive layers, the neuron wiggle can be used to recover neuron signs individually. Consider a change in input $\Delta_{k}$ sufficiently small to stay within the linear region but which maximises $\hat{\eta_{k}}(x^{*}+\Delta_{k})$. The general intuition is that the ReLU blocks the change caused by $\Delta_{k}$ on the inactive side of $\eta$. Therefore if our sign guess is correct, the output should display a large change. If our sign guess is incorrect, we shouldn’t see a particular change. We can observe a correct change only if we can sufficiently maximise $\hat{\eta_{k}}(x^{*}+\Delta_{k})$ without maximising at the same time other neurons. We refer the reader to the original article [6] for further details. This makes the neuron wiggle a probabilistic method which depends on many factors, notably the architecture of the network. As a consequence, the neuron wiggle is ran on many critical points $x^{*}$ of $\eta_{k}$ and the sign is chosen according to the overall results, with a certain confidence level $\alpha$. $\alpha=1$ indicates that all points agree, whereas $\alpha=0$ implies an even split. Canales-Martínez et al. report very convincing results on a $3072-256^{(8)}-10$ network, recovering the sign of all but ten neurons in this very large network. The neuron wiggle is further scrutinised in [12], where the authors recommend running an exhaustive search on the signs of the neurons with confidence level $\alpha<0.75$.
@@ 2.4 Overview of Existing Sign-Extraction Methods
Figure 2: Left: Original signature extraction from [8]. Right: Proposed improvements. Two error-inducing steps in the original attack are coloured on the left. Improvements match the colour of the step they address. Further precision improvements for signature extraction are marked on the right.
3 Improving Signature Extraction
To the best of our knowledge, all prior work has been able to apply signature extraction on very shallow neural networks only (3 layers maximum). Specifically, according to Table 1 in [8], the deepest model they attempted to attack had an architecture of 40201010140-20-10-10-1, consisting of only $3$ hidden layers. On the other hand, Table 2 in [12] shows that they attempted to attack a model with an architecture of $784-16^{(8)}-1$, which had $8$ hidden layers. However, despite running the attack for over $36$ hours, they failed to recover the fourth hidden layer. Although [6] demonstrated that their sign extraction was effective across all layers of a complex $3072-256^{(8)}-10$ model, they supposed in their experiments that the signatures had been at this stage correctly extracted, without running the two processes together.
Motivated by these limitations, we analyze the signature extraction method and identify two challenges preventing its successful application on deeper layers. We then propose solutions to overcome them. The left part of Fig. 2 illustrates the signature extraction workflow and highlights where these challenges arise.
(a) Original signature extraction.
(b) After discarding deeper points.
(c) After signature intersections.
(d) After discarding deeper components.
Figure 3: Gradual improvement of signature extraction results on layer 4 of Model II with $3,000$ critical points (due to space constraints, only the largest 16 components are displayed). Each component on the target layer is labelled with its associated neuron on top. (a) Original signature extraction from [8]. (b) After discarding deeper critical points (see Section 3.2), deeper components either disappear or have a smaller size. (c) After intersecting critical points with insufficient ranks (see Section 3.1), components with rank deficiency disappear. (d) After discarding deeper components (see Section 3.2), all remaining components are in the target layer. The only unrecovered component corresponds to an always-off neuron $\eta^{(4)}_{5}$.
We provide now a brief overview of each failure. First, the signature extraction relies on solving a system of equations to recover the weights of the target neuron associated with all active neurons in the previous layer. However, as we go deeper into the network, the rank of this system might not match the number of active neurons on the previous layer. This mismatch can lead to an underdetermined system, resulting in incorrect signature extraction. Such components are in red in Fig. 3.
Second, the component selection of the largest $d_{i}$ components of size greater than 2 is not a good approach. Indeed, our experiments show that this strategy is only valid for shallow networks where there are comparatively fewer critical points on deeper layers. The right part of Fig. 2 gives an outline of our improvements. Components from deeper layers are in brown in Fig. 3.
Our methods for resolving the issues of incorrect signature extraction and deeper components are given respectively in Section 3.1 and Section 3.2. We show in Fig. 3, as an illustration, how each improvement we propose allows us to gradually reach a correct extraction the 4th layer of Model II with architecture $784-8^{(8)}-1$. The same approach naturally extends to subsequent layers.
Moreover, we present in Section 3.3 several numerical precision‑improvement strategies that increase extraction efficiency, and provide additional precision refinements for wide layers in the context of an end-to-end attack. The rightmost part of Fig. 2 indicates when they are applied.
3.1 Increasing Rank with Subspace Intersections
$\cdots\cdots\cdots\cdotsInput$ $x$
Figure 4: Solving for a neuron’s weights from certain critical points can yield an underdetermined system. The target neuron is shown in black; active neurons are in blue. At input $x$, $\operatorname{rank}(\Gamma^{(1)}_{x})=2$, so $\operatorname{rank}(\Gamma^{(2)}_{x})\leq 2$ even though three neurons are active in layer 22; consequently, the layer-2 system has no unique solution.
The issue: insufficient rank. When extracting the signature, our goal is to recover a partial signature $y_{x}\in\mathbb{R}^{d_{i-1}}$ of a neuron $\eta^{(i)}$ at its critical point $x$. Consequently, the solution space should be of dimension 1. To determine $y_{x}$, we solve $(\Gamma_{x}^{(i-1)}\Delta)\cdot y_{x}=\partial^{2}_{\Delta}f(x)$ using a sufficient number of random directions $\Delta\in\mathbb{R}^{d_{0}}$. The resulting partial row $y_{x}$ consists of the coefficients for the neurons in layer i1i-1 that are active at input $x$, entries on inactive indices are zero. We denote by $s_{x}^{(i-1)}$ the number of active neurons in layer i1i-1 and by $r^{(i-1)}_{x}$ the rank of the map $\Gamma^{(i-1)}_{x}\in\mathbb{R}^{d_{i-1}\times d_{0}}$. If $r_{x}^{(i-1)}<s_{x}^{(i-1)}$, the system is underdetermined and $y_{x}$ cannot be uniquely determined. This typically occurs when some earlier layer k<i1k<i-1 has fewer active neurons than layer i1i-1 (see Fig. 4), since $\Gamma_{x}^{(i-1)}=I_{x}^{(i-1)}A^{(i-1)}\cdots I_{x}^{(1)}A^{(1)}$ and thus:
$\text{rank}(\Gamma_{x}^{(i-1)})\leq\min(\{\text{rank}(I_{x}^{(k)})\}_{1\leq k\leq i-1})\neq\text{rank}(I_{x}^{(i-1)})\text{ in general.}$
If $X$ denotes the random variable representing the number of active neurons per layer, then $\mathbb{P}(r_{x}^{(i-1)}<s_{x}^{(i-1)})=1-\mathbb{P}(X\geq s_{x}^{(i-1)})^{i-1}$. Because of the exponent (i1)(i-1), this (failure) probability approaches $1$ rapidly with depth, making extraction beyond the first few layers increasingly unlikely.
Signature intersections. If the rank is insufficient, we no longer have a unique solution but rather a space of solutions. For a critical point $x$ in the layer $i$, the solution space is given by $\mathcal{S}_{x}=L_{x}+\ker(\Gamma_{x}^{(i-1)})$, where $L_{x}$ is any particular solution and $\ker(\cdot)$ denotes the kernel. Suppose we have another critical point $x^{\prime}$ with solution space $\mathcal{S}_{x^{\prime}}=L_{x^{\prime}}+\ker(\Gamma_{x^{\prime}}^{(i-1)}).$ We want to intersect $\mathcal{S}_{x}$ and $\mathcal{S}_{x^{\prime}}$ into a space of smaller dimension containing $A^{(i)}_{k}$. Doing this for multiple critical points should give us a space of dimension $1$, yielding $A^{(i)}_{k}$ up to a scaling factor.
Let’s study this in more details. Since we recover rows up to a scalar factor, we seek to compute the intersection of $\mathcal{S}_{x}$ and $\lambda\mathcal{S}_{x^{\prime}}$, where $\lambda$ is a scalar factor equal to the ratio of the row scalar factors of $x$ and $x^{\prime}$. Let $(e_{1},\ldots,e_{k})$ be a basis of $\ker(\Gamma_{x}^{(i-1)})$, and let $(f_{1},\ldots,f_{m})$ be a basis of $\ker(\Gamma_{x^{\prime}}^{(i-1)})$. We then solve the following system in $\mathbb{R}^{d_{i-1}}$:
$L_{x}+\mu_{1}e_{1}+\dots+\mu_{k}e_{k}=\lambda L_{x^{\prime}}+\lambda_{1}f_{1}+\dots+\lambda_{m}f_{m},$
where the unknowns are $\mu_{1},\dots,\mu_{k}$ and $\lambda,\lambda_{1},\dots,\lambda_{m}\in\mathbb{R}$. This system consists of $d_{i-1}$ equations in $\mathbb{R}$. However, only the equations corresponding to the active neurons in the layer i1i-1 for both inputs $x$ and $x^{\prime}$ are relevant. To ensure that we do not merge spaces from critical points associated with different neurons, we overdetermine the system. Indeed, an overdetermined system with random coefficients is inconsistent with very high probability, and thus a system of $\ell>N$ equations with $N$ unknowns typically doesn’t have a solution except if there is some ground truth behind it. Therefore, we impose the following merging condition:
$\ell>1+|\ker(\Gamma_{x}^{(i-1)})|+|\ker(\Gamma_{x^{\prime}}^{(i-1)})|=1+k+m,$
where $\ell$ is the number of relevant equations.
Intersections allow us to exploit at least 90% of critical points on the target layer when attacking for example the Model II. By comparison, the original attack gives a uniquely solvable system (and thus a correct signature) for only about $6\%$ of layer-4 critical points, with even worse results in deeper layers. We could intersect subspaces three by three to have a lower condition for merging, allowing for the use of more critical points. However, this would considerably slow down the attack. For this reason, all our results use pairwise merges.
We provide in Fig. 3(c) an experiment showing how intersecting the critical-point solution spaces prunes components affected by rank deficiency, making the correct ones stand out.
3.2 Deeper Layers’ Influence on Extraction